# SciEx: Benchmarking Large Language Models on Scientific Exams with Human Expert Grading and Automatic Grading

### v1

@@ Abstract
With the rapid development of Large Language Models (LLMs), it is crucial to have benchmarks which can evaluate the ability of LLMs on different domains. One common use of LLMs is performing tasks on scientific topics, such as writing algorithms, querying databases or giving mathematical proofs. Inspired by the way university students are evaluated on such tasks, in this paper, we propose SciEx - a benchmark consisting of university computer science exam questions, to evaluate LLMs ability on solving scientific tasks. SciEx is (1) multilingual, containing both English and German exams, and (2) multi-modal, containing questions that involve images, and (3) contains various types of freeform questions with different difficulty levels, due to the nature of university exams. We evaluate the performance of various state-of-the-art LLMs on our new benchmark. Since SciEx questions are freeform, it is not straightforward to evaluate LLM performance. Therefore, we provide human expert grading of the LLM outputs on SciEx. We show that the free-form exams in SciEx remain challenging for the current LLMs, where the best LLM only achieves 59.4% exam grade on average. We also provide detailed comparisons between LLM performance and student performance on SciEx. To enable future evaluation of new LLMs, we propose using LLM-as-a-judge to grade the LLM answers on SciEx. Our experiments show that, although they do not perform perfectly on solving the exams, LLMs are decent as graders, achieving 0.948 Pearson correlation with expert grading.
@@
Tu Anh Dinh, Carlos Mullov, Leonard Bärmann, Zhaolin Li, Danni Liu, Simon Reiß, Jueun Lee, Nathan Lerzer, Fabian Ternava, Jianfeng Gao, Alexander Waibel, Tamim Asfour, Michael Beigl, Rainer Stiefelhagen, Carsten Dachsbacher, Klemens Böhm, Jan Niehues Karlsruhe Institute of Technology, Karlsruhe, Germany {firstname}.{lastname}@kit.edu
@@ 1 Introduction
In this paper, we introduce a new benchmark, termed SciEx (Scientific Exams), designed to evaluate this capability. Inspired by the way students are evaluated in university, we created the benchmark by evaluating the performance of LLMs on university computer science exams. SciEx’s questions are in various formats, from multiple choice to open-ended, thus suitable to evaluate LLM’s capabilities of generating free-text answers that fit the requirements of the questions. It is multilingual, containing exams in both German and English. It is multimodal, as exam questions can also contain figures. The set of questions is a good mix of different difficulty levels since they are designed for university exams. This enables us to evaluate LLMs on different levels, and we found that stronger LLMs tend to perform better on more difficult questions.
@@
  • SciEx [^1] We release SciEx under CC BY-NCSA 4.0 license. Code: https://github.com/TuAnh23/SciEx. Data: https://huggingface.co/datasets/tuanh23/SciEx. - a freeform, multimodal, multilingual benchmark consisting of university computer science exams, outputs of various LLMs on the exams, and expert grading of the LLM output.
@@ 2 Related Work
Other studies have instead provided scientific benchmarks with open-ended questions. Some examples are Qasper Dasigi et al. (2021) and ARC-DA Bhakthavatsalam et al. (2021). However, Qasper only focuses on questions about NLP papers rather than on general computer science topics. ARC-DA is closer to our work, since it contains open-ended questions taken from science exam and quiz sources. However, these are created by converting questions that were originally multiple-choice, thus not covering certain types of typical freeform questions (e.g. those that require mathematical proofs, or long explanations).
@@
One recent approach is to use LLMs for evaluation, termed “LLM-as-a-judge”. Liu et al. (2023); Chiang and Lee (2023); Zheng et al. (2024) find that, although still prone to biases, LLM-as-a-judge for textual modality has high agreement with human scoring when a strong judge LLM is used. However, when including images, Chen et al. (2024) find that the performance of LLM-as-a-judge is no longer as well correlated to human judgment. Nevertheless, LLM-as-a-judge is a promising way to perform scalable evaluation.
@@ 3 The SciEx Benchmark
There are in total 154 unique questions. Each question is annotated with (1) the maximum points that can be achieved and (2) a difficulty level among Easy, Medium, Hard. Most questions are provided with gold reference answer and average student performance. The detailed per-question statistics are shown in Table 1.
@@ 3.1 Data Creation
The data creation process is described as follows.
@@
We then give the LLM answers back to the lecturers, who proceed with grading the LLM output the same way they would grade student answers. We anonymized the LLMs’ names in order to avoid bias during exam grading. We also build a user interface for collecting the grades (see Appendix C for more details).
@@ 3.2 Automatic Grading
In addition to expert grading, we also provide automatic grading using LLM-as-a-judge, so that we can evaluate future LLMs on SciEx. We use the stronger models, i.e., Mixtral, Llama3 and GPT-4V, to conduct the grading. Given a tuple containing question, answer, and maximum score, we ask the LLMs to output a single score between 0 and the maximum. We include reference answers to the grading prompt. We ask the LLMs to provide a chain-of-thought reasoning Wei et al. (2022) before giving the grade. We also include examples for grading in the prompt, so-called few-shot judge Zheng et al. (2024). Each example is a tuple consisting of a question, an answer, and the grade. We try out different settings to select the examples:
  • Same question: select examples from the same question but with an answer-grade pair for a different LLM examinee. This mimics the real-life scenario where we use the expert resource to grade some answers of the same exam, then use it to guide the LLM graders.
  • Same exam: select examples from a different question but the same exam. Here the examples are in the same domain as the question-answer pair in consideration. This mimics the real-life scenario where, e.g., we have expert grading on exam of the same course from previous years to guide LLM graders.
  • Different exam: select examples from a different question and a different exam. This mimics the real-life scenario where, e.g., we have expert grading for an exam of another course to guide the LLM graders.
Intuitively, the example-selection settings above have decreasing levels of relevance to the actual grading query, but increasing easiness to collect.
@@ 4.1.1 General Observations
Grade (%) German Scale
@@ 4.1.2 Influential Factors
Figure 1(a) shows the influence of the difficulty level on the examinee grades. As can be seen, the student performance aligns with the difficulty level of the questions: they perform better on easier questions. Some weaker LLMs, e.g., Mixtral, Qwen, GPT-3.5, Llava, align with the students. However, the stronger LLMs, i.e., Claude and GPT-4V, perform better on harder questions. This is an indication that, difficulty levels from human perspective do not always align with LLMs’ perspective. This is also confirmed by looking at the correlations between the LLM grades and the student grades. These correlations are between 0.4 to 0.6, indicating that LLM grades and the student grades are not highly correlated.
In Figure 1(b), we plot the difference between LLM scores and student scores. The stronger LLMs, i.e., Claude and GPT-4V, outperform the students the most on hard questions. Weaker LLMs, on the other hand, generally fall behind students the most on hard questions. A potential reason is that math-type easy questions are hard for the LLMs, while long-text hard questions are easy for them.
Looking at each difficulty level independently, we observe that the ranking of the LLMs changes across different levels. This aligns with the findings made by Li et al. (2024), where they show that the LLM rankings change on a subset of evaluation prompts that are artificially labeled as hard.
@@
Figure 2 shows the influence of images on the difference between LLM and student scores. Trivially, the LLMs that cannot handle images perform poorly on the image-related questions. The strong, multi-modal LLMs, i.e., Claude and GPT4, outperform the students on both image-related questions and text-only questions, but the performance gap is still larger for text-only questions. Llava, although can handle images, still falls behind student performance by a large margin on image-related questions. This shows that LLMs’ image-handling capability is still not as advanced as for text.
@@
Since some models are not made to deal with images, or with languages other than English, we additionally analyze LLMs’ performance on text-only and English-only questions. On this subset of questions, the grades obtained by the models are generally better, and more models would outperform the student average. More details can be found in Appendix D.
@@ 4.2 Qualitative Analysis
The graders observed some common behaviors made by the LLMs. Some solutions of the LLMs were good language-wise but low-quality content-wise. For students, good language usually correlates strongly with good content. The LLMs tend to output lengthy answers, since, unlike the students, they do not have a time constraint. Some LLMs even ignore when the question specifies that they should ”answer briefly“. There are also some failure cases: (1) Claude refuses to answer the question with "I apologize, but I do not feel comfortable providing answers related to …" or (2) some LLMs get stuck in decoding loops. Sometimes, instead of answering the question, LLMs give some text that is (or seems) related to the task; rephrase the task; or describe how a task of this nature may be approached in general.
@@
The LLMs tend to fail on the math-related questions, even the basic ones. For example, they miscount the number of words in a piece of text, or have trouble comparing numbers. For questions that require writing mathematical proof in the TGI exam, all LLMs except for GPT-4V and Claude failed. For GPT-4V and Claude, they are able to pass the TGI exam. Their mistakes are more in line with those that students would make. That is, they are often not successful when making actual proof, and the points where the proof breaks sometimes are the same as the students. Even the better models handle simple geometry questions poorly and/or struggle to follow instructions of a simple algorithm.
@@
Table 4: LLM grading’s Pearson correlation to expert grading on the question level. Note that there are only a single scores for zero-shot, since they do not have different shot settings.
Reasoning Ability
Image Handling
4.3 Automatic Grading
In this section, we evaluate the performance of LLM-as-a-judge approach to automatic grading. We use the expert grades as the gold standard to evaluate automatic graders. We use Pearson correlation on the normalized scores as our metric. Since the LLMs are asked to provide the scores on the same scale as the expert scores, we also provide the Root Mean Squared Error (RMSE) on the originally-scaled scores as a secondary metric. Note that RMSE would correctly put more weight on the questions that have more points, however, it is not as easily interpretable as the Pearson correlation. Therefore, we only report RMSE in Appendix E.2. The main results are discussed as follows.
4.3.1 General
@@
On the exam level, LLM-as-a-judge performs well for automatic grading. The best Pearson correlation to expert grading on the exam level, at 0.948, is achieved by GPT-4V. The open-source Llama3 achieves 0.883 Pearson correlation to expert grading. These high correlations indicate that, although being far from perfect in solving SciEx exams (discussed in Section 4.1.1), the stronger LLMs are quite reliable for grading the exams. This is useful since we would have to rely less on expert grading to evaluate newly-developed LLMs’ performance on SciEx. The details of graders’ performance under different settings on the exam level is in Appendix E.1.
On the sentence level, the performance of LLM-as-a-judge is shown in Table 4. The highest Pearson correlation to expert grading achieved by the LLMs is now around 0.7, which is lower than on the exam level, but still quite high. Surprisingly, the performance of GPT-4V on grading image-related questions is quite comparable to grading text-only questions. This contradicts to the finding made by Chen et al. (2024). This could potentially be due to the small number of image-related questions in SciEx, thus the results might not be generalizable.
@@
The performance of the graders on the question level is shown in Table 4. We observe that adding more examples (shots) and adding the reference answers in the prompt generally increases the performance of the LLM graders. GPT-4V is the strongest grader, followed by Llama3 and Mixtral. This shows that proprietary LLMs are still stronger as judge, aligning with previous studies Zheng et al. (2024).
@@ 4.3.2 Grader-specific Behaviours
As can be seen from Table 4, Mixtral has the worst performance on grading the exams. We observe that Mixtral tends to give full points to the answers. Without reference and without examples (0-shot), the portion of answers where Mixtral outputs full points is 67.6%, significantly higher than Llama3 and GPT-4V, at 19.1% and 15.1%, respectively. As a result, Mixtral’s precision on giving full points, at 0.181, is much lower than Llama3 and GPT-4V, at 0.380 and 0.527 respectively. As we add more examples and/or add the reference answer to the prompt, the problem is lessened. More details can be found in Appendix E.3.

### v3

@@ Abstract
With the rapid development of Large Language Models (LLMs), it is crucial to have benchmarks which can evaluate the ability of LLMs on different domains. One common use of LLMs is performing tasks on scientific topics, such as writing algorithms, querying databases or giving mathematical proofs. Inspired by the way university students are evaluated on such tasks, in this paper, we propose SciEx - a benchmark consisting of university computer science exam questions, to evaluate LLMs’ ability on solving scientific tasks. SciEx is (1) multilingual, containing both English and German exams, and (2) multi-modal, containing questions that involve images, and (3) contains various types of freeform questions with different difficulty levels, due to the nature of university exams. We evaluate the performance of various state-of-the-art LLMs on our new benchmark. Since SciEx questions are freeform, it is not straightforward to evaluate LLM performance. Therefore, we provide human expert grading of the LLM outputs on SciEx. We show that the free-form exams in SciEx remain challenging for the current LLMs, where the best LLM only achieves 59.4% exam grade on average. We also provide detailed comparisons between LLM performance and student performance on SciEx. To enable future evaluation of new LLMs, we propose using LLM-as-a-judge to grade the LLM answers on SciEx. Our experiments show that, although they do not perform perfectly on solving the exams, LLMs are decent as graders, achieving 0.948 Pearson correlation with expert grading.
@@
Tu Anh Dinh, Carlos Mullov, Leonard Bärmann, Zhaolin Li, Danni Liu, Simon Reiß, Jueun Lee, Nathan Lerzer, Jianfeng Gao, Fabian Ternava, Tobias Röddiger, Alexander Waibel, Tamim Asfour, Michael Beigl, Rainer Stiefelhagen, Carsten Dachsbacher, Klemens Böhm, Jan Niehues Karlsruhe Institute of Technology, Karlsruhe, Germany {firstname}.{lastname}@kit.edu
@@ 1 Introduction
In this paper, we introduce a new benchmark, termed SciEx (Scientific Exams), designed to evaluate this capability. Inspired by the way students are evaluated in university, we created the benchmark by evaluating the performance of LLMs on university computer science exams. SciEx’s questions are in various formats, from multiple choice to open-ended, thus making it suitable to evaluate LLM’s capabilities of generating free-text answers that fit the requirements of the questions. It is multilingual, containing exams in both German and English. It is multimodal, as exam questions can also contain figures. The set of questions is a good mix of different difficulty levels since they are designed for university exams. This enables us to evaluate LLMs on different levels, and we found that stronger LLMs tend to perform better on more difficult questions.
@@
  • SciEx [^1] We release SciEx under CC BY-NC-SA 4.0 license. Code: https://github.com/TuAnh23/SciEx. Data: https://huggingface.co/datasets/tuanh23/SciEx. - a freeform, multimodal, multilingual benchmark consisting of university computer science exams, outputs of various LLMs on the exams, and expert grading of the LLM output.
@@ 2 Related Work
Other studies have instead provided scientific benchmarks with open-ended questions. Some examples are Qasper Dasigi et al. (2021) and ARC-DA Bhakthavatsalam et al. (2021). However, Qasper only focuses on questions about NLP papers rather than on general computer science topics. ARC-DA is closer to our work, since it contains open-ended questions taken from science exams and quiz sources. However, these are created by converting questions that were originally multiple-choice, thus not covering certain types of typical freeform questions (e.g. those that require mathematical proofs, or long explanations).
@@
One recent approach is to use LLMs for evaluation, termed “LLM-as-a-judge”. Liu et al. (2023); Chiang and Lee (2023a); Zheng et al. (2024) find that, although still prone to biases, LLM-as-a-judge for textual modality has high agreement with human scoring when a strong judge LLM is used. However, when including images, Chen et al. (2024) find that the performance of LLM-as-a-judge is no longer as well correlated to human judgment. Nevertheless, LLM-as-a-judge is a promising way to perform scalable evaluation.
@@ 3 The SciEx Benchmark
There are in total 154 unique questions. Each question is annotated with (1) the maximum points that can be achieved and (2) a difficulty level among Easy, Medium, Hard. Most questions are provided with gold reference answers and average student performance. The detailed per-question statistics are shown in Table 1.
@@ 3.1 Data Creation
The data creation process is as follows.
@@
The detailed prompts for LLMs to generate the answers are shown in Appendix C.
@@
We then give the LLM answers back to the lecturers, who proceed with grading the LLM output the same way they would grade student answers. We anonymized the LLMs’ names in order to avoid bias during exam grading. We also build a user interface for collecting the grades (see Appendix E for more details).
@@ 3.2 Automatic Grading
In addition to expert grading, we also provide automatic grading using LLM-as-a-judge, so that we can evaluate future LLMs on SciEx. We use the stronger models, i.e., Mixtral, Llama3 and GPT-4V, to conduct the grading. Given a tuple containing question, answer, and maximum score, we ask the LLMs to output a single score between 0 and the maximum. We include reference answers to the grading prompt. We ask the LLMs to provide chain-of-thought reasoning Wei et al. (2022); Chiang and Lee (2023b) before giving the grade. We also include examples for grading in the prompt, so-called few-shot judge Zheng et al. (2024). Each example is a tuple consisting of a question, an answer, and the expert-provided grade. We try out different settings to select the examples, as described below.
Let’s say we want to grade Question M from Exam A, answered by Examinee X. Then the shot examples can be chosen in one of the three ways:
  • Same question: Select examples from the same Question M from Exam A, but answered by a different Examinee Y. This mimics the real-life scenario where we use the expert resource to grade some answers of the same exam, then use it to guide the LLM graders.
  • Same exam: select examples from a different Question N from the same Exam A, Answered by a different Examinee Y. Here the examples are in the same domain as the question-answer pair in consideration. This mimics the real-life scenario where, e.g., we have expert grading on exams of the same course from previous years to guide LLM graders.
  • Different exam: select examples from a different Question N, from a different Exam B, Answered by a different Examinee Y. This mimics the real-life scenario where, e.g., we have expert grading for an exam of another course to guide the LLM graders.
Examinee Y and Question N are chosen randomly. For Exam B, we opt to select the exams that do not heavily require images for simplicity in the prompt.
Intuitively, the example-selection settings above have decreasing levels of relevance to the actual grading query, but increasing easiness to collect. The detailed prompts for LLMs to grade answers are shown in Appendix D.
@@ 4.1 Quantitative Analysis
We also report on the German grade scale. In the German scale, the grades range from $1.01.0$ to $5.05.0$, where $1.01.0$ is the highest grade and $4.04.0$ is the passing threshold. The detailed mapping from the scores to the German grade scale is defined by the lectures, adjusted based on the overall performance of the students taking the exams.
We compare the performance of the LLMs to students from different aspects: language, difficulty level, and modality, i.e., questions with or without images.
@@ 4.1.1 General Observations
Grade (%) $\uparrow$ German Scale $\downarrow$
@@ 4.1.2 Influential Factors
Figure 1(a) shows the influence of the difficulty level on the examinee grades. As can be seen, the student performance aligns with the difficulty level of the questions: they perform better on easier questions. Some weaker LLMs, e.g., Mixtral, Qwen, GPT-3.5, Llava, align with the students. However, the stronger LLMs, i.e., Claude and GPT-4V, perform better on harder questions. This is an indication that difficulty levels from human perspective do not always align with LLMs’ perspective. This is also confirmed by looking at the Pearson correlations between the LLMs’ grades and the student average grades on the question level. These correlations are between 0.4 and 0.6, indicating that LLM grades and the student grades are not highly correlated.
One possible explanation for the mismatch between LLMs performance and question difficulty level could be that, in some exams, there can be some “template questions”, i.e., questions that are repeated over the years, where students can just learn by heart how to systematically solve them. While this would be marked as “easy” by the lecturer, it might not be as easy for the LLMs, since the LLMs are not previously exposed to these “template questions”. Another potential explanation is that math-type easy questions are hard for the LLMs, while long-text hard questions are easy for them.
In Figure 1(b), we plot the difference between LLM scores and student scores. The stronger LLMs, i.e., Claude and GPT-4V, outperform the students the most on hard questions. Weaker LLMs, on the other hand, generally fall behind students the most on hard questions. Looking at each difficulty level independently, we observe that the ranking of the LLMs changes across different levels. This aligns with the findings made by Li et al. (2024), where they show that the LLM rankings change on a subset of evaluation prompts that are artificially labeled as hard.
@@
Figure 2 shows the influence of images on the difference between LLM and student scores. Recall that for the text-only LLMs, we exclude the images and only pass the question text to the models. Trivially, the text-only LLMs perform poorly on the image-related questions. The strong, multi-modal LLMs, i.e., Claude and GPT4, outperform the students on both image-related questions and text-only questions, but the performance gap is still larger for text-only questions. Llava, although can handle images, still falls behind student performance by a large margin on image-related questions. This shows that LLMs’ image-handling capability is still not as advanced as for text.
@@
Since some models are not made to deal with images, or with languages other than English, we additionally analyze LLMs’ performance on text-only and English-only questions. On this subset of questions, the grades obtained by the models are generally better, and more models would outperform the student average. More details can be found in Appendix F.
@@ 4.2 Qualitative Analysis
The graders observed some common behaviors made by the LLMs. Some solutions of the LLMs were good language-wise but low-quality content-wise. For students, good language usually correlates strongly with good content. The LLMs tend to output lengthy answers, since, unlike the students, they do not have a time constraint. Some LLMs even ignore when the question specifies that they should ”answer briefly“. There are also some failure cases, although not frequent: (1) Claude refuses to answer the question with "I apologize, but I do not feel comfortable providing answers related to …" or (2) some LLMs get stuck in decoding loops. Sometimes, instead of answering the question, LLMs give some text that is (or seems) related to the task; rephrase the task; or describe how a task of this nature may be approached in general.
@@
The LLMs tend to fail on the math-related questions, even the basic ones. For example, they miscount the number of words in a piece of text, or have trouble comparing numbers. For questions that require writing mathematical proof in the TGI exam, all LLMs except for GPT-4V and Claude failed. For GPT-4V and Claude, they are able to pass the TGI exam. Their mistakes are more in line with those that students would make. That is, they are often not successful when making actual proof, and the points where the proof breaks sometimes are the same as the students. Even the better models handle simple geometry questions poorly and/or struggle to follow the instructions of a simple algorithm.
Reasoning Ability
Image Handling
4.3 Automatic Grading
In this section, we evaluate the performance of LLM-as-a-judge approach to automatic grading. We use the expert grades as the gold standard to evaluate automatic graders. We use Pearson correlation on the normalized scores as our metric. Since the LLMs are asked to provide the scores on the same scale as the expert scores, we also provide the Root Mean Squared Error (RMSE) on the originally-scaled scores as a secondary metric. Note that RMSE would correctly put more weight on the questions that have more points, however, it is not as easily interpretable as the Pearson correlation. Therefore, we only report RMSE in Appendix G.2. The main results are discussed as follows.
4.3.1 General
@@
Table 4: LLM grading’s Pearson correlation to expert grading on the question level. Note that there are only single scores for zero-shot, since they do not have different shot settings.
@@
On the exam level, LLM-as-a-judge performs well for automatic grading. The best Pearson correlation to expert grading on the exam level, at 0.948, is achieved by GPT-4V. The open-source Llama3 achieves 0.883 Pearson correlation to expert grading.
The LLM ranking based on average exam-level grades provided by the GPT-4V grader in comparison to expert grading is shown in Table 5. As can be seen, the ranking is quite identical, except for Mixtral and Qwen’s positions being swapped.
Expert grader GPT-4V grader
Examinee Avg. grade Examinee Avg. grade
(%, sorted) (%, sorted)
Claude 59.4 Claude 57.7
GPT-4V 58.2 GPT-4V 56.2
Mixtral 41.1 Qwen 42.0
Qwen 35.4 Mixtral 38.2
GPT-3.5 32.8 GPT-3.5 38.0
Mistral 25.9 Mistral 24.6
Llava 21.5 Llava 24.2
Table 5: LLM examinees ranking with expert grader and GPT-4V grader.
The high correlations between expert grading and LLM-as-a-judge grading indicate that, although being far from perfect in solving SciEx exams (discussed in Section 4.1.1), the stronger LLMs are quite reliable for grading the exams. This is useful since we would have to rely less on expert grading to evaluate newly developed LLMs’ performance on SciEx. The details of graders’ performance under different settings on the exam level are in Appendix G.1.
On the question level, the performance of LLM-as-a-judge is shown in Table 4. The highest Pearson correlation to expert grading achieved by the LLMs is now around 0.7, which is lower than on the exam level, but still quite high. Surprisingly, the performance of GPT-4V on grading image-related questions is quite comparable to grading text-only questions. This contradicts the finding made by Chen et al. (2024). This could potentially be due to the small number of image-related questions in SciEx, thus the results might not be generalizable.
@@
The performance of the graders on the question level is shown in Table 4. We observe that adding examples (shots) and adding reference answers in the prompt generally increases the performance of the LLM graders. GPT-4V is the strongest grader, followed by Llama3 and Mixtral. This shows that proprietary LLMs are still stronger as judges, aligning with previous studies Zheng et al. (2024).
@@ 4.3.2 Grader-specific Behaviours
As can be seen from Table 4, Mixtral has the worst performance on grading the exams. We observe that Mixtral tends to give full points to the answers. Without reference and without examples (0-shot), the portion of answers where Mixtral outputs full points is 67.6%, significantly higher than Llama3 and GPT-4V, at 19.1% and 15.1%, respectively. As a result, Mixtral’s precision on giving full points, at 0.181, is much lower than Llama3 and GPT-4V, at 0.380 and 0.527 respectively. As we add more examples and/or add the reference answer to the prompt, the problem is lessened. More details can be found in Appendix G.3.